OpenStudy (anonymous):

Solve the equation –5 × 1 × 11 × 4 step by step. List each step, along with the property you used during that step

OpenStudy (anonymous):

-5 x 1 = -5 -5 x 11 = -55 -55 x 4 = -220

OpenStudy (anonymous):

What properties did you use to get the answer?

OpenStudy (anonymous):

I don't know what that means. I just multiplied the numbers together.

Parth (parthkohli):

I don't find it to be an “equation”.

OpenStudy (anonymous):

A. Commutative property B. Associative property C. Identity property

OpenStudy (anonymous):

Associative property I guess as I did (((-5x1)x11)x4)

OpenStudy (anonymous):

Also, the identity property for -5x1

OpenStudy (anonymous):

As that is the property that states all numbers multiplied by 1 stay the same.

OpenStudy (anonymous):

Thank you so much!!!!

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

Feel free to give me a medal.

OpenStudy (anonymous):

Thank you :D